# Qingdao 266237, China

### v1

@@ 3.2 Multi-party Private Set Union
  • Parameters. $m$ parties $P_{1},\cdots,P_{m}$, where $P_{\textnormal{ld}}$ is the leader, $\textnormal{ld}\in[m]$. Size $n$ of input sets. The bit length $l$ of set elements.
  • Functionality. On input $X_{i}=\{x_{i}^{1},\cdots,x_{i}^{n}\}\subseteq\{0,1\}^{l}$ from $P_{i}$, give $\bigcup_{i=1}^{m}X_{i}$ output to $P_{\textnormal{ld}}$.
@@
3.3 Batch Oblivious Pseudorandom Function
Oblivious pseudorandom function (OPRF) [FIPR05] is a central primitive in the area of PSO. Kolesnikov et al. [KKRT16] introduced batched OPRF, which provides a batch of OPRF instances in the following way. In the $i$th instance, the receiver $\mathcal{R}$ has an input $x_{i}$; the sender $\mathcal{S}$ learns a PRF key $k_{i}$ and $\mathcal{R}$ learns $\mathsf{PRF}(k_{i},x_{i})$. Note that $\mathcal{R}$ evaluates the PRF on only one point per key. The batch OPRF functionality is described formally in Figure 3.
  • Parameters. Sender $\mathcal{S}$. Receiver $\mathcal{R}$. Batch size $B$. The bit length $l$ of set elements. Some PRF family $\mathsf{PRF}:\{0,1\}^{*}\times\{0,1\}^{l}\to\{0,1\}^{\gamma}$.
  • Functionality. On input $\vec{x}\subseteq(\{0,1\}^{l})^{B}$ from $\mathcal{R}$,
    • For each $i\in[B]$, choose uniform PRF key $k_{i}$.
    • For each $i\in[B]$, define $f_{i}=\mathsf{PRF}(k_{i},x_{i})$.
    • Give vector $\vec{k}=(k_{1},\cdots,k_{B})$ to $\mathcal{S}$ and vector $\vec{f}=(f_{1},\cdots,f_{B})$ to $\mathcal{R}$.
Figure 3: Batch OPRF Functionality $\mathcal{F}_{\mathsf{bOPRF}}$
3.4 Oblivious Key-Value Stores
A key-value store [PRTY20, GPR+21, RR22, BPSY23] is a data structure that compactly represents a desired mapping from a set of keys to corresponding values.
Definition 1
A key-value store is parameterized by a set $\mathcal{K}$ of keys, a set $\mathcal{V}$ of values, and a set of functions $\mathsf{H}$, and consists of two algorithms:
  • $\mathsf{Encode_{H}}$ takes as input a set of key-value pairs $\{(k_{i},v_{i})|i\in[n]\}$ and outputs an object $D$ (or, with statistically small probability, an error indicator $\perp$).
  • $\mathsf{Decode_{H}}$ takes as input an object $D$, a key $k$, and outputs a value $v$.
Correctness. For all $A\subseteq\mathcal{K}\times\mathcal{V}$ with distinct keys:
$(k,v)\in A\ \ and\ \perp\neq D\leftarrow\mathsf{Encode_{H}}(A)\Longrightarrow\mathsf{Decode_{H}}(D,k)=v$
Obliviousness. For all distinct $\{k_{1}^{0},\cdots,k_{n}^{0}\}$ and all distinct $\{k_{1}^{1},\cdots,k_{n}^{1}\}$, if $\mathsf{Encode_{H}}$ does not output $\perp$ for $\{k_{1}^{0},\cdots,k_{n}^{0}\}$ or $\{k_{1}^{1},\cdots,k_{n}^{1}\}$, then the distribution of $D_{0}\leftarrow\mathsf{Encode_{H}}(\{(k_{1}^{0},v_{1}),\cdots,(k_{n}^{0},v_{n})\})$ is computationally indistinguishable to $D_{1}\leftarrow\mathsf{Encode_{H}}(\{(k_{1}^{1},v_{1}),\cdots,(k_{n}^{1},v_{n})\})$, where $v_{i}\leftarrow\mathcal{V}$ for $i\in[n]$.
A KVS is an Oblivious KVS (OKVS) if it satisfies the obliviousness property. we also require an additional randomness property [ZCL+23] from the OKVS.
Randomness. For any $A=\{(k_{1},v_{1}),\cdots,(k_{n},v_{n})\}$ and $k^{*}\notin\{k_{1},\cdots,k_{n}\}$, the output of $\mathsf{Decode_{H}}(D,k^{*})$ is statistically indistinguishable to that of uniform distribution over $\mathcal{V}$, where $D\leftarrow\mathsf{Encode_{H}}(A)$.
@@ 3.5 Batch Oblivious Programmable Pseudorandom Function
Oblivious programmable pseudorandom function (OPPRF) [KMP+17, PSTY19, CGS22, RS21, RR22] is an extension of OPRF, which lets $\mathcal{S}$ program a PRF $F$ so that it has specific random outputs for some specific inputs and pseudorandom outputs for all other inputs. This kind of PRF with the additional property that on a certain programmed set of inputs the function outputs programmed values is called programmable PRF (PPRF) [PSTY19]. $\mathcal{R}$ evaluates the OPPRF with no knowledge of whether it learns a programmed output of $F$ or just a pseudorandom value. The batch OPPRF functionality is given in Figure 4.
  • Parameters. Sender $\mathcal{S}$. Receiver $\mathcal{R}$. Batch size $B$. The bit length $l$ of keys. The bit length $\gamma$ of values. An OKVS scheme $(\mathsf{Encode_{H}},\mathsf{Decode_{H}})$.
@@
    • The value sets $V_{1},\cdots,V_{B}$, where $\lvert K_{i}\rvert=\lvert V_{i}\rvert$ for every $i\in[B]$, and $V_{i}(j)\in\{0,1\}^{\gamma}$ for every $i\in[B]$ and $1\leq j\leq\lvert K_{i}\rvert$.
@@
    • Choose uniformly random and independent PPRF key $k_{i}$, for each $i\in[B]$;
@@
3.6 Hashing to Bin
The hashing to bin technique was introduced by Pinkas et al. [PSZ14, PSSZ15], which is originally applied to construct two-party PSI protocol. At the high level, the receiver $\mathcal{R}$ uses hash functions $h_{1},h_{2},h_{3}$ to assign its items to $B$ bins via Cuckoo hashing [PR04], so that each bin has at most one item. The hashing process uses eviction and the choice of which of the bins is used depends on the entire set. Using the same hash functions and simple hashing, sender $\mathcal{S}$ assigns each of its items $x$ to all of the bins $h_{1}(x),h_{2}(x),h_{3}(x)$, so that if the item $x$ is also in $\mathcal{R}$’s set and is mapped into the bin $b\in\{h_{1}(x),h_{2}(x),h_{3}(x)\}$ by Cuckoo hashing, then the bin $b$ of $\mathcal{S}$’s simple hash table certainly contains $x$.
We write simple hashing with the following notation:
@@ 3.6 Hashing to Bin
This expression means to hash the items of $X$ into $B$ bins using simple hashing with hash functions $h_{1},h_{2},h_{3}:\{0,1\}^{*}\to[B]$. The output is the simple hash table denoted by $\mathcal{T}^{1},\cdots,\mathcal{T}^{B}$, where for each $x\in X$ we have $\mathcal{T}^{h_{i}(x)}=\{x\|i|i=1,2,3\}$. [^7] Appending the index of the hash function is helpful for dealing with edge cases like $h_{1}(x)=h_{2}(x)=i$, which happen with non-negligible probability. Without appending $i$, the OKVS would contain the identical value $F(k_{i},x)$, which leaks the fact that a collision $h_{1}(x)=h_{2}(x)$ occurred. Such an event is input-dependent so cannot be simulated.
We write Cuckoo hashing with the following notation:
@@
This expression means to hash the items of $X$ into $B$ bins using Cuckoo hashing on hash functions $h_{1},h_{2},h_{3}:\{0,1\}^{*}\to[B]$. The output is the Cuckoo hash table denoted by $\mathcal{C}^{1},\cdots,\mathcal{C}^{B}$, where for each $x\in X$ there is some $i\in\{1,2,3\}$ such that $\mathcal{C}^{h_{i}(x)}=x\|i$. Some Cuckoo hash positions do not matter, corresponding to empty bins. We use these symbols throughout the subsequent sections.
@@ 3.7 Secret-Shared Private Equality Test
Secret-shared private equality test protocol (ssPEQT) [PSTY19, CGS22] can be viewed as an extreme case of ssPMT when the sender $\mathcal{S}$’s input set has only one item. In Figure 5 we formally define this functionality.
  • Parameters. Two parties $P_{1},P_{2}$. The bit length $\gamma$ of inputs.
  • Functionality. On input $x$ from $P_{1}$ and input $y$ from $P_{2}$, sample two random bits $a,b$ under the constraint that $a\oplus b=\bigwedge_{i=1}^{\gamma}(1\oplus x[i]\oplus y[i])$. Namely, if $x=y$, $a\oplus b=1$. Otherwise $a\oplus b=0$. Give $a$ to $P_{1}$ and $b$ to $P_{2}$.
@@ 3.9 Multi-Party Secret-Shared Shuffle
Multi-party secret-shared shuffle has the capability to permute the share vectors of all parties in a random permutation unknown to any coalition of $m-1$ parties. Then it refreshs all shares. The functionality is given in Figure 7.
Eskandarian et al. [EB22] propose an online-efficient protocol by extending [CGP20] to the multi-party setting. In the offline phase, each party generates a random permutation and a set of correlated vectors called share correlation. In the online phase, each party permutes and refreshes the share vectors efficiently using share correlation. We give the functionality of share correlation and details of the online protocol in Appendix 0.C.
@@
3.10 Multi-Key Rerandomizable Public-Key Encryption
Gao et al. [GNT23] introduce the concept of multi-key rerandomizable public-key encryption (MKR-PKE), which is a variant of PKE with several additional properties. Let $\mathcal{SK}$ denote the space of secret keys which forms an abelian group under the operation $+$, $\mathcal{PK}$ the space of public keys which forms an abelian group under the operation $\cdot$, $\mathcal{M}$ the space of plaintexts, and $\mathcal{C}$ the space of ciphertexts. MKR-PKE is a tuple of PPT algorithms $(\mathsf{Gen},\mathsf{Enc},\mathsf{ParDec},\mathsf{Dec},\mathsf{ReRand})$ such that:
@@ 3.10 Multi-Key Rerandomizable Public-Key Encryption
MKR-PKE requires the following additional properties besides those of PKE with indistinguishable multiple encryptions. (cf. Appendix 0.B):
3.10.1 Partially Decryptable
For any two pairs of keys $(sk_{1},pk_{1})\leftarrow\mathsf{Gen}(1^{\lambda})$ and $(sk_{2},pk_{2})\leftarrow\mathsf{Gen}(1^{\lambda})$ and any $x\in\mathcal{M}$, it holds that
@@ 3.10.1 Partially Decryptable
3.10.2 Rerandomizable
For any $pk\in\mathcal{PK}$ and any $x\in\mathcal{M}$, it holds that
@@ 3.10.2 Rerandomizable
[GNT23] uses elliptic curve based ElGamal encryption [ElG85] to instantiate MKR-PKE. The concrete EC MKR-PKE is described in Appendix 0.D.
@@
We note that only a few of elliptic curves support efficient encoding from bit-strings to EC points. Therefore, the plaintext space of EC MKR-PKE is generally restricted to EC points to guarantee rerandomizable property.
@@ 4 Batch Secret-Shared Private Membership Test
The batch secret-shared private membership test (batch ssPMT) is a central building block in both our SK-MPSU and PK-MPSU protocols. In this section, we formally introduce this functionality and provide a construction built by the aforementioned primitives.
The batch ssPMT is a two-party protocol implementing a batch of ssPMT instances between a sender and a receiver. Assuming a batch size of $B$, the sender $\mathcal{S}$ inputs $B$ sets $X_{1},\cdots,X_{B}$, while the receiver $\mathcal{R}$ inputs $B$ elements $x_{1},\cdots,x_{B}$. Consequently, $\mathcal{S}$ and $\mathcal{R}$ receive secret shares of a bit vector of size $B$, whose $i$th bit is 1 if $x_{i}\in X_{i}$, otherwise 0. The batch ssPMT functionality is presented in Figure 8 and the construction is given in Figure 9.
@@
  • Functionality. On inputs $X_{1},\cdots,X_{B}$ from $\mathcal{S}$ and input $\vec{x}$ from $\mathcal{R}$, for each $i\in[B]$, sample two random bits $e_{S}^{i},e_{R}^{i}$ under the constraint that if $x_{i}\in X_{i},e_{S}^{i}\oplus e_{R}^{i}=1$, otherwise $e_{S}^{i}\oplus e_{R}^{i}=0$. Give $\vec{e}_{S}=(e_{S}^{1},\cdots,e_{S}^{B})$ to $\mathcal{S}$ and $\vec{e}_{R}=(e_{R}^{1},\cdots,e_{R}^{B})$ to $\mathcal{R}$.
@@
      The parties invoke $B$ instances of $\mathcal{F}_{\mathsf{ssPEQT}}$, where in the $i$th instance $\mathcal{S}$ inputs $s_{i}$ and $\mathcal{R}$ inputs $t_{i}$. In the end, $\mathcal{S}$ receives $e_{S}^{i}\in\{0,1\}$ and $\mathcal{R}$ receives $e_{R}^{i}\in\{0,1\}$.
@@
  • Correctness. According to the functionality of batch OPPRF, if $x_{i}\in X_{i}$, then $\mathcal{R}$ receives $t_{i}=s_{i}$. Subsequently, in the $i$th instance of ssPEQT, they input $s_{i}$ and $t_{i}$ which are equal, hence their outputs satisfy $e_{S}^{i}\oplus e_{R}^{i}=1$. Conversely, if $x_{i}\notin X_{i}$, then $\mathcal{R}$ receives a pseudorandom value $t_{i}$. The probability of $t_{i}=s_{i}$ equals $2^{-\gamma}$ and any $t_{i}=s_{i},i\in[B]$ occurs with a probability of $B\cdot 2^{-\gamma}$. As we set $\gamma\geq\sigma+\log B$, so $B\cdot 2^{-\gamma}\leq 2^{-\sigma}$, which means the probability of any $t_{i}=s_{i}$ occurring is negligible. After the invocation of ssPEQT, we conclude that if $x_{i}\notin X_{i}$, $e_{S}^{i}$ and $e_{R}^{i}$ satisfy $e_{S}^{i}\oplus e_{R}^{i}=0$ with overwhelming probability.
@@
Protocol $\Pi_{\mathsf{bssPMT}}$ securely realizes $\mathcal{F}_{\mathsf{bssPMT}}$ in the $(\mathcal{F}_{\mathsf{bOPRF}},\mathcal{F}_{\mathsf{ssPEQT}})$-hybrid model.
The security of the protocol follows immediately from the security of the batch OPPRF and the ssPEQT functionalities.
  • Complexity Analysis. We set $B=O(n)$ to be consistent with our MPSU protocols. One thing to clarify is that the asymptotic complexity, to a great extent, depends on the choices of the instantiation of batch OPRF and the generation method of Beaver Triples [Bea91] in ssPEQT. In consideration of concrete efficiency, we employ subfield vector oblivious linear evaluation (subfield-VOLE) [BCG+19a, BCG+19b, RRT23] for both options. A comprehensive complexity analysis is in Appendix 0.F.1. We summarize the total costs as below:
    • Offline phase. The computation complexity of each party is $O(\gamma n\log n)$. The communication complexity of each party is $O(t\lambda\log(\gamma n/t))$. The round complexity is $O(1)$.
    • Online phase. The computation complexity of each party is $O(n)$. The communication complexity of each party is $O(\gamma n)$. The round complexity is $O(\log\gamma)$.

### v4

@@ 3.2 Multi-party Private Set Union
    • Parameters. $m$ parties $P_{1},\cdots,P_{m}$, where $P_{1}$ is the leader. Size $n$ of input sets. The bit length $l$ of set elements.
    • Functionality. On input $X_{i}=\{x_{i}^{1},\cdots,x_{i}^{n}\}\subseteq\{0,1\}^{l}$ from $P_{i}$ ($i\in[m]$), output $\bigcup_{i=1}^{m}X_{i}$ to $P_{1}$.
@@ 3.3 Batch Oblivious Programmable Pseudorandom Function
    Oblivious pseudorandom function (OPRF) [FIPR05] is a central primitive in the area of PSO. Kolesnikov et al. [KKRT16] introduced batched OPRF, which provides a batch of OPRF instances. In the $i$th instance, the sender $\mathcal{S}$ learns a PRF key $k_{i}$, while the receiver $\mathcal{R}$ inputs $x_{i}$ and learns $\mathsf{PRF}(k_{i},x_{i})$.
    Oblivious programmable pseudorandom function (OPPRF) [KMP+17, PSTY19, CGS22, RS21, RR22] is an extension of OPRF, which lets $\mathcal{S}$ program a PRF $F$ so that it has specific uniform outputs for some specific inputs and pseudorandom outputs for all other inputs. This kind of PRF with the additional property that the function outputs programmed values on a certain programmed set of inputs is called programmable PRF (PPRF) [PSTY19]. $\mathcal{R}$ evaluates the OPPRF with no knowledge of whether it learns a programmed output of $F$ or just a pseudorandom value. The batch OPPRF functionality is given in Figure 3.
    • Parameters. Sender $\mathcal{S}$. Receiver $\mathcal{R}$. Batch size $B$. The bit length $l$ of keys. The bit length $\gamma$ of values.
@@
      • The value sets $V_{1},\cdots,V_{B}$, where $\lvert K_{i}\rvert=\lvert V_{i}\rvert$, $i\in[B]$.
@@
      • Choose a uniform PPRF key $k_{i}$, for $i\in[B]$;
@@
    3.4 Hashing to Bins
    The hashing to bins technique was introduced by Pinkas et al. [PSZ14, PSSZ15] to construct two-party PSI. At a high level, the receiver $\mathcal{R}$ uses hash functions $h_{1},h_{2},h_{3}$ to assign its items to $B$ bins via Cuckoo hashing [PR04], so that each bin has at most one item. The Cuckoo hashing process uses eviction and the choice of bins for each item depends on the entire set. On the other hand, the sender $\mathcal{S}$ assigns each of its items $x$ to all bins $h_{1}(x),h_{2}(x),h_{3}(x)$ via simple hashing. This guarantees that for each item $x$ of $\mathcal{R}$, if $x$ is mapped into the $b$th bin of Cuckoo hash table ($b\in\{h_{1}(x),h_{2}(x),h_{3}(x)\}$), and $x$ is in $\mathcal{S}$’s set, then the $b$th of simple hash table certainly contains $x$.
    We denote simple hashing with the following notation:
@@ 3.4 Hashing to Bins
    This expression represents hashing the items of $X$ into $B$ bins using simple hashing with hash functions $h_{1},h_{2},h_{3}:\{0,1\}^{*}\to[B]$. The output is a simple hash table denoted by $\mathcal{T}^{1},\cdots,\mathcal{T}^{B}$, where for each $x\in X$ we have $\mathcal{T}^{h_{i}(x)}\supseteq\{x\|i|i=1,2,3\}$. [^8] Appending the index of the hash function is helpful for dealing with edge cases like $h_{1}(x)=h_{2}(x)=i$, which happen with non-negligible probability.
    We denote Cuckoo hashing with the following notation:
@@
    This expression represents hashing the items of $X$ into $B$ bins using Cuckoo hashing with hash functions $h_{1},h_{2},h_{3}:\{0,1\}^{*}\to[B]$. The output is a Cuckoo hash table denoted by $\mathcal{C}^{1},\cdots,\mathcal{C}^{B}$, where for each $x\in X$ there is some $i\in\{1,2,3\}$ such that $\mathcal{C}^{h_{i}(x)}=\{x\|i\}$. Some Cuckoo hash positions are irrelevant, corresponding to empty bins. We use these symbols throughout the subsequent sections.
@@ 3.5 Secret-Shared Private Equality Test
    Secret-shared private equality test protocol (ssPEQT) [PSTY19, CGS22] can be viewed as an extreme case of ssPMT when the sender $\mathcal{S}$’s input set size is 1. Figure 4 defines its functionality.
    • Parameters. Two parties $P_{1},P_{2}$. The input bit length $\gamma$.
    • Functionality. On input $x$ from $P_{1}$ and input $y$ from $P_{2}$, sample two random bits $a,b$ s.t. if $x=y$, $a\oplus b=1$. Otherwise $a\oplus b=0$. Give $a$ to $P_{1}$ and $b$ to $P_{2}$.
@@ 3.7 Multi-Party Secret-Shared Shuffle
    Multi-party secret-shared shuffle functionality works by randomly permuting the share vectors of all parties and then refreshing all shares, ensuring that the permutation remains unknown to any coalition of $m-1$ parties. The formal functionality is given in Figure 6.
    Eskandarian et al. [EB22] proposed an online-efficient protocol by extending [CGP20] to the multi-party setting. In the offline phase, each party generates a random permutation and a set of correlated vectors called share correlations. In the online phase, each party efficiently permutes and refreshes the share vectors using these share correlations.
@@
    Gao et al. [GNT23] introduced the concept of multi-key rerandomizable public-key encryption (MKR-PKE), a variant of PKE with several additional properties. Let $\mathcal{SK}$ denote the space of secret keys, which forms an abelian group under the operation $+$, and $\mathcal{PK}$ denote the space of public keys, which forms an abelian group under the operation $\cdot$. $\mathcal{M}$ denotes the space of plaintexts, and $\mathcal{C}$ denotes the space of ciphertexts. MKR-PKE is a tuple of PPT algorithms $(\mathsf{Gen},\mathsf{Enc},\mathsf{ParDec},\mathsf{Dec},\mathsf{ReRand})$ such that:
@@
    MKR-PKE is an IND-CPA secure PKE scheme that requires the following additional properties:
    • Partially Decryptable. For any two pairs of keys $(sk_{1},pk_{1})\leftarrow\mathsf{Gen}(1^{\lambda}),(sk_{2},pk_{2})\leftarrow\mathsf{Gen}(1^{\lambda})$ and any $x\in\mathcal{M}$,
@@
    • Rerandomizable. For any $pk\in\mathcal{PK}$ and any $x\in\mathcal{M}$,
@@
      Gao et al. [GNT23] make use of elliptic curve (EC) based ElGamal encryption [ElG85] to instantiate MKR-PKE. The concrete EC MKR-PKE is described in Appendix 0.C.
@@
      To the best of our knowledge, there is few elliptic curves that support efficient encoding and decoding between bit-strings and EC points. Therefore, the plaintext space of EC MKR-PKE is generally restricted to EC points to guarantee rerandomizable property.
@@ 4 Batch Secret-Shared Private Membership Test
    The batch secret-shared private membership test (batch ssPMT) is a central building block in our SK-MPSU and PK-MPSU protocols. It is a two-party protocol that implements multiple instances of ssPMT between a sender $\mathcal{S}$ and a receiver $\mathcal{R}$. Given a batch size of $B$, $\mathcal{S}$ inputs $B$ sets $X_{1},\cdots,X_{B}$, while $\mathcal{R}$ inputs $B$ elements $x_{1},\cdots,x_{B}$. As a result, $\mathcal{S}$ and $\mathcal{R}$ receive secret shares of a bit vector of size $B$, where the $i$th bit is 1 if $x_{i}\in X_{i}$, 0 otherwise. The batch ssPMT functionality is presented in Figure 7 and the construction is given in Figure 8, built with batch OPPRF and ssPEQT.
@@
    • Functionality. On inputs $X_{1},\cdots,X_{B}$ from $\mathcal{S}$ and input $\vec{x}$ from $\mathcal{R}$, for each $i\in[B]$, sample two random bits $e_{S}^{i},e_{R}^{i}$ s.t. if $x_{i}\in X_{i},e_{S}^{i}\oplus e_{R}^{i}=1$, otherwise $e_{S}^{i}\oplus e_{R}^{i}=0$. Give $\vec{e}_{S}=(e_{S}^{1},\cdots,e_{S}^{B})$ to $\mathcal{S}$ and $\vec{e}_{R}=(e_{R}^{1},\cdots,e_{R}^{B})$ to $\mathcal{R}$.
@@
        The parties invoke $B$ instances of $\mathcal{F}_{\mathsf{ssPEQT}}$, where in the $i$th instance $\mathcal{S}$ inputs $s_{i}$ and $\mathcal{R}$ inputs $t_{i}$. In the end, $\mathcal{S}$ and $\mathcal{R}$ receives $e_{S}^{i},e_{R}^{i}\in\{0,1\}$ respectively.
@@
    Protocol $\Pi_{\mathsf{bssPMT}}$ securely realizes $\mathcal{F}_{\mathsf{bssPMT}}$ in the $(\mathcal{F}_{\mathsf{bOPPRF}},\mathcal{F}_{\mathsf{ssPEQT}})$-hybrid model.
    • Correctness. According to the batch OPPRF functionality, if $x_{i}\in X_{i}$, $\mathcal{R}$ receives $t_{i}=s_{i}$. Then in the $i$th instance of ssPEQT, since $s_{i}=t_{i}$, $e_{S}^{i}\oplus e_{R}^{i}=1$. Conversely, if $x_{i}\notin X_{i}$, $\mathcal{R}$ receives a pseudorandom value $t_{i}$. The probability that any $t_{i}=s_{i}$ for $i\in[B]$ is $B\cdot 2^{-\gamma}$. By setting $\gamma\geq\sigma+\log B$, we have $B\cdot 2^{-\gamma}\leq 2^{-\sigma}$, meaning the probability of any such collision is negligible [^9] In MPSU, the batch OPPRF is invoked multiple times, and the lower bound of $\gamma$ is relevant to the total number of batch OPPRF invocations. Refer to  0.E.2 for more details.. After the invocation of ssPEQT, we have that if $x_{i}\notin X_{i}$, $e_{S}^{i}\oplus e_{R}^{i}=0$ with overwhelming probability.
    • Security. The security of the protocol follows immediately from the security of batch OPPRF and ssPEQT functionalities.
    In our MPSU protocols, the batch ssPMT is always combined with the hashing to bins technique. The combined construction has good efficiency, together with linear computation and communication in term of $n$, which mainly benefits from the following technical advances: First, we follow the paradigm in [PSTY19] to construct batch OPPRF from batch OPRF and oblivious key-value store (OKVS) [PRTY20, GPR+21, RR22, BPSY23]. By leveraging the technique to amortize communication, the total communication of computing $B=O(n)$ instances of OPPRF is equal to the total number of items $3n$. Second, we utilize subfield vector oblivious linear evaluation (subfield-VOLE) [BCG+19a, BCG+19b, RRT23] to instantiate batch OPRF and the construction in [RR22] to instantiate OKVS. This ensures the computation complexity of batch OPPRF of size $O(n)$ to scale linearly with $n$. A comprehensive analysis is in Appendix 0.E.2.